# A solution of the problem of standard compact Clifford-Klein forms

### v1

@@ Abstract
We solve the long standing problem of classification of standard compact Clifford-Klein forms of homogeneous spaces of simple non-compact real Lie groups. The result is that standard compact Clifford-Klein forms always arise from triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ of real Lie algebras such that $\mathfrak{h}\subset\mathfrak{g},\mathfrak{l}\subset\mathfrak{g}$, $\mathfrak{g}$ is simple and absolutely simple, $\mathfrak{h},\mathfrak{l}$ are (non-compact) reductive, $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}$, and the intersection $\mathfrak{h}\cap\mathfrak{l}$ is compact. The consequence of this is the following characterization of proper co-compact actions of reductive Lie subgroups $L\subset G$ on a homogeneous spaces $G/H$ determined by absolutely simple real Lie group $G$ and a closed reductive subgroup $H$: $L$ acts on $G/H$ properly and co-compactly if and only if $G=H\cdot L$ and $H\cap L$ is compact.
Key words and phrases: proper action, homogeneous space, semisimple Lie algebra. 2020 Mathematics Subject Classification: 17B05, 17B10, 53C30, 57S25, 57S30.
@@ 1 Introduction
A homogeneous space $G/H$ admits a standard compact Clifford-Klein form, if there exists a closed reductive subgroup $L\subset G$ such that $L$ acts properly and co-compactly on $G/H.$
This notion was first introduced in [24]. A list of such spaces was given in [23] and [25]. The term "standard" was first used in [19]. Standard Clifford-Klein forms are determined by triples $(G,H,L)$, where $G$ is a semisimple real linear Lie group, and $H$ and $L$ are reductive subgroups. The general theory [24] shows that we can restrict ourselves to considering semisimple Lie subgroups $H,L$ (see Section 2.1 for details), and we will do that throughout this article.
@@
The proof of this result is well known: all triples in Table 1 are Lie algebra decompositions in the sense of Onishchik: $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}$ [29]. It follows from this work, that $G=H\cdot L$ as well, and that $L$ acts transitively on $G/H$. One obtains a natural diffeomorphism $G/H\cong L/L\cap H$, and so if $H\cap L$ is compact, $L$ acts properly and co-compactly on $G/H$. It is important to note that these are the only known examples of compact standard Clifford-Klein forms (with $\mathfrak{g}$ simple and $\mathfrak{h}$ proper and non-compact).
In this paper we prove that this Table contains all non-trivial triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ with semisimple $\mathfrak{h}$ and $\mathfrak{l}$ generating standard compact Clifford-Klein forms.
@@
Assume that $\mathfrak{g}$ is absolutely simple and $\mathfrak{h},$ $\mathfrak{l}$ are proper semisimple subalgebras of non-compact type. The only Lie algebra triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ which yield compact standard Clifford-Klein forms are contained in Table 1.
The discussion in Section 2 shows that Theorem 2 yields a complete classification of all standard compact Clifford-Klein forms determined by triples $(G,H,L)$ with $G$ simple and $H,L$ reductive. Note that if $\mathfrak{g}$ is simple exceptional then, by [4], for any proper reductive subalgebra $\mathfrak{h}\subset\mathfrak{g}$ of non-compact type the corresponding homogeneous space does not admit compact standard Clifford-Klein forms. Therefore, in the proof of Theorem 2 we may assume that $\mathfrak{g}$ is classical simple.
It is also important to stress that the proof of Theorem 2 yields an interesting characterization of proper co-compact actions of reductive Lie groups on homogeneous spaces $G/H$ determined by simple absolutely simple Lie groups $G$ and (non-compact) reductive subgroups.
@@
Let $G/H$ be a homogeneous space determined by a simple absolutely simple Lie group $G$ and a reductive (non-compact) closed subgroup $H$. A reductive Lie subgroup $L\subset G$ acts on $G/H$ properlyand co-compactly, if and only if $G=H\cdot L$ and $H\cap L$ is compact.
There are various reasons for studying standard compact Clifford-Klein forms. One of them comes from spectral analysis on pseudo-Riemannian manifolds [19], [20]. In this area, one is interested in constructing quotients $M_{\Gamma}=\Gamma\backslash M$ of discrete groups of isometries of a pseudo-Riemannian homogeneous manifold $M=G/H$. However, it may happen that $M$ does not admit infinite discontinuous groups of isometries (the Calabi-Markus phenomenon). On the other hand, there exist examples of manifolds $G/H$ admitting large discontinuous groups such that $M_{\Gamma}$ is compact, or of finite volume. In this case, we will call $M_{\Gamma}$ a compact Clifford-Klein form, or a Clifford-Klein form of finite volume. Thus, it is a challenging problem to understand when do such compact quotients exist. Suppose a Lie subgroup $L\subset G$ acts properly on $G/H$. Then the action of any discrete subgroup $\Gamma$ of $L$ is automatically properly discontinuous, and this action is free whenever $\Gamma$ is torsion-free. Moreover, if $\Gamma$ is co-compact in $L$, and $L$ acts co-compactly on $G/H$, then the quotient $\Gamma\backslash G/H$ is compact. A quotient $\Gamma\backslash G/H$ is called standard if $\Gamma$ is contained in a reductive subgroup $L$ acting properly on $G/H$. Thus, the first question which arises is to find standard quotients, which is equivalent to finding triples $(G,H,L)$ as in Definition 1. Partial results in this direction were obtained in [4] and [32].
@@
One of the most challenging problems in this area of research is a problem of T. Kobayashi: does any homogeneous space $G/H$ admitting compact Clifford-Klein form admits also a standard one? Clearly, one of the ways to attack it is to understand all standard Clifford-Klein forms.
In a more geometric vein, if $\Gamma\subset G$ acts freely and co-compactly on a homogeneous space $G/H$ with non-compact $G$ and $H$, any $G$-invariant geometric structure descends onto the quotient, which yields compact manifolds with such structures. The topology of such manifolds can be understood, and therefore, these manifolds can serve as testing examples. For instance, we mention [3].
@@
In the context of this work, one should mention that there are many non-existence results for compact Clifford-Klein forms (not necessarily standard). For example, there are no standard Clifford-Klein forms of exceptional Lie groups [4]. Topological obstructions to compact Clifford-Klein forms were found in [27],[28], [33] together with examples. A procedure of constructing homogeneous spaces with no compact Clifford-Klein forms was obtained in [5]. On the positive side, all compact standard Clifford-Klein of symmetric spaces were classified in [32]. The existence of non-standard compact Clifford-Klein forms is a deep mathematical problem of a different nature. We refer to a very recent work [26]. Acknowledgement. We thank T. N. Venkataramana for valuable discussions during the meeting "Zariski dense subgroups, number theory and geometric applications" held at ICTS, Bangalore in January 2024, and for pointing out an interpretation of our results described by Theorem 3.
2 Notation and summary of [6]
@@ 2.1 Properties of Lie algebra embeddings yielding compact Clifford-Klein forms
Assume that $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ is a standard triple, which means that $\mathfrak{g}$ is absolutely simple, $\mathfrak{h},\mathfrak{l}$ are reductive subalgebras of $\mathfrak{g}$ and $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ induces a compact standard Clifford-Klein form. Throughout we work with root systems of complex semisimple Lie algebras and with real root systems of their real forms. Also, we use some relations between them, which we reproduce here in order to fix notation (see [30], Section 4.1). We may assume that there exists a Cartan involution $\theta$ of $\mathfrak{g}$ such that $\theta(\mathfrak{h})=\mathfrak{h},$ $\theta(\mathfrak{l})=\mathfrak{l}.$ Hence the restriction of $\theta$ is a Cartan involution of $\mathfrak{h}$ and $\mathfrak{l}.$ We obtain the following Cartan decompositions
@@
$\mathfrak{g}^{c}=\mathfrak{j}^{c}+\sum_{\alpha_{c}\in\Delta}\mathfrak{g}_{\alpha_{c}}.$
The Lie algebra $\mathfrak{j}:=i\mathfrak{t}+\mathfrak{a}$ is a real form of $\mathfrak{j}^{c}$ with
$\mathfrak{j}=\{A\in\mathfrak{j}^{c}\ |\ \alpha_{c}(A)\in\mathbb{R}\ \textrm{for any}\ \alpha_{c}\in\Delta\}.$
The weights of the adjoint representation of $\mathfrak{a}$ constitute a root system $\Sigma\subset\mathfrak{a}^{*}$ (not necessarily reduced). There is a natural map $pr:\Delta\rightarrow\Sigma$ given by $\alpha_{c}\rightarrow\alpha_{c}|_{\mathfrak{a}}$. It is easy to see that $pr(\Delta)=\Sigma\cup\{0\}.$ Therefore, we will call $pr(\alpha_{c})=\alpha_{c}|_{\mathfrak{a}}\in\mathfrak{a}^{*}$ (if non-zero) the restricted roots. Thus, $\Sigma=\{\alpha:=\alpha_{c}|_{\mathfrak{a}}\ |\ \alpha_{c}\in\Delta\}\setminus\{0\},$
$\mathfrak{g}_{\alpha}:=\sum_{\begin{subarray}{c}\alpha_{c}\in\Delta\\ \alpha_{c}|_{\mathfrak{a}}=\alpha\end{subarray}}\mathfrak{g}_{\alpha_{c}}\cap\ \mathfrak{g}.$
The reader should keep in mind, that we use the following convention. Roots of complex Lie algebra $\mathfrak{g}^{c}$ are denoted by $\alpha_{c},\beta_{c},,...$, the corresponding restrictions are $\alpha,\beta,...$ and so on. Vectors from the root spaces $\mathfrak{g}_{\alpha_{c}}$ of $\mathfrak{g}^{c}$ and $\mathfrak{g}_{\alpha}$ from $\mathfrak{g}$ are always denoted by $x_{\alpha},x_{\beta},...$. The meaning of these will be always clear from the context. However, to avoid clumsy notation, in several cases we write $\alpha_{i}$ for complex roots. In each case we state explicitly, which root (complex or its restriction) is used.
@@
$\mathfrak{g}=\mathfrak{m}_{0}+\mathfrak{a}+\sum_{\alpha\in\Sigma}\mathfrak{g}_{\alpha}.$
Let $B$ be the Killing form of $\mathfrak{g}.$ It extends onto $\mathfrak{g}^{c}$ as a Killing form of $\mathfrak{g}^{c}$ and we denote it by the same letter. Consider the compact real form $\mathfrak{g}_{u}:=\mathfrak{k}+i\mathfrak{p},$ of $\mathfrak{g}^{c},$ and denote by $\tau$ the complex conjugation in $\mathfrak{g}^{c}$ with respect to $\mathfrak{g}_{u}.$ Also we denote by $\sigma$ the complex conjugation of $\mathfrak{g}^{c}$ with respect to $\mathfrak{g}$. It is well known ([31], p. 228) that there is a Hermitian non-degenerate form on $\mathfrak{g}^{c}$ given by the formula $B_{\tau}(X,Y)=-B(X,\tau(Y)).$ Choose a set of positive restricted roots $\Sigma^{+}\subset\Sigma.$ There is a system of positive roots $\Delta^{+}\subset\Delta$ such that
$\Sigma^{+}=\{\alpha_{c}|_{\mathfrak{a}}\ |\ \alpha_{c}\in\Delta^{+}\}-\{0\}.$
@@
as well as $\mathfrak{n}_{h}^{-},\mathfrak{n}_{l}^{-}$ for $\mathfrak{h}$ and $\mathfrak{l}.$ Since $\theta(\mathfrak{g}_{\alpha})=\mathfrak{g}_{-\alpha}$ for all $\alpha\in\Sigma$ ([17], Proposition 5.9) we get $\theta(\mathfrak{n})=\mathfrak{n}^{-}$.

### v3

@@ Abstract
We solve the problem of classification of standard compact Clifford-Klein forms of homogeneous spaces of simple non-compact real Lie groups under the following assumptions:
  1. 1.
    either standard compact Clifford-Klein forms are given by triples$(G,H,L),$ with $G,H,L$ linear connected which arise from triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ of real Lie algebras with $\mathfrak{g}$ absolutely simple, and at least one of the complexified subalgebras $\mathfrak{h}^{c}\subset\mathfrak{g}^{c}$ or $,\mathfrak{l}^{c}\subset\mathfrak{g}^{c}$, is regular, or
  2. 2.
    $\mathfrak{g},\mathfrak{h},\mathfrak{l}$ are absolutely simple.
The solution is that such triples yield standard compact Clifford-Klein forms if and only if $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}$ and $\mathfrak{h}\cap\mathfrak{l}$ is compact.
Key words and phrases: proper action, homogeneous space, semisimple Lie algebra, Clifford-Klein form, Lie algebra decomposition. 2020 Mathematics Subject Classification: 17B05, 17B10, 53C30, 57S25, 57S30.
@@ 1.1 The problem and motivation
A homogeneous space $G/H$ admits a standard compact Clifford-Klein form, if there exists a discrete subgroup $\Gamma\subset G$ and closed reductive subgroup $L\subset G$ containing $\Gamma$ as a co-compact lattice such that $\Gamma$ (and thus $L$) acts properly and co-compactly on $G/H.$
This notion was first introduced in [24]. A list of such spaces was given in [23] and [25]. The term "standard" was first used in [19]. Standard Clifford-Klein forms are determined by triples $(G,H,L)$, where $H$ and $L$ are reductive subgroups. The general theory [24] shows that we can restrict ourselves to considering semisimple Lie subgroups $H,L$ (see Section 2.2 for details), and we will do that throughout this article. The following is well known.
@@
It is important to note that these are the only known examples of homogeneous spaces admitting compact Clifford-Klein forms (with $\mathfrak{g}$ simple and $\mathfrak{h}$ proper and non-compact). All Lie algebra triples in Table 1 come from [32], Table 2. In this work, Onishchik classifies all triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ such that $\mathfrak{g}$ is simple, $\mathfrak{h},\mathfrak{l}$ are reductive and $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}$. Such triples are called Lie algebra decompositions. We omit the complete description of the embeddings $\mathfrak{h}\hookrightarrow\mathfrak{g}$, referring to [32]. Foe example, in the last four lines one of the subalgebras is embedded via an appropriate spin representation (see [32], the proof of Theorem 4.1). Also, the last three lines in the table are already contained in the table as special cases with $n=1$. However, we prefer to leave it in the form given by Table 2 in [32], for the convenience of the reader. Let us also note that according to the general theory of compact Clifford-Klein forms [24], it is allowed to interchange the roles of $\mathfrak{h}$ and $\mathfrak{l}$, and we will do that throughout this article.
In this paper we prove that under some mild restrictions Table 1 exhausts all the possibilities for compact standard Clifford-Klein forms. The main result is as follows.
@@
Assume that $\mathfrak{g}$ is absolutely simple, and $\mathfrak{h},\mathfrak{l}$ satisfy the following:
  1. 1.
    either the complexified embedding $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$ is regular, or
  2. 2.
    both $\mathfrak{h}$ and $\mathfrak{l}$ are proper absolutely simple subalgebras of non-compact type in $\mathfrak{g}$.
Then only Lie algebra triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ which yield compact standard Clifford-Klein forms are Lie algebra decompositions, that is $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}$.
Note that if $\mathfrak{g}$ is absolutely simple exceptional then, by [6], for any proper reductive subalgebra $\mathfrak{h}\subset\mathfrak{g}$ of non-compact type the corresponding homogeneous space does not admit compact standard Clifford-Klein forms. Therefore, in the proof of Theorem 2 we may assume that $\mathfrak{g}$ is classical simple.
It is interesting that the proof of Theorem 2 yields a characterization of proper co-compact actions of reductive Lie groups on homogeneous spaces $G/H$ determined by simple absolutely simple Lie groups $G$ and (non-compact) absolutely simple subgroups (we thank T. N. Venkataramana for this observation).
@@
Let $G/H$ be a homogeneous space determined by an absolutely simple Lie groups $G,H,$ $H\subset G$. An absolutely simple Lie subgroup $L\subset G$ acts on $G/H$ properly and co-compactly, if and only if $G=H\cdot L$ and $H\cap L$ is compact.
There are various reasons for studying compact Clifford-Klein forms (we mention only some of them):
  1. 1.
    spectral analysis on pseudo-Riemannian manifolds [19], [20],[27],
  2. 2.
@@
    a problem of T. Kobayashi: does any homogeneous space $G/H$ admitting compact Clifford-Klein form admits also a standard one? Clearly, one of the ways to attack it is to understand all standard Clifford-Klein forms,
  3. 3.
    constructing various compact manifolds with particular geometric structures. For example, a homogeneous space of reductive type is an important example of a homogeneous pseudo-Riemannian manifold. This observation was exploited in [4] to explore Einstein metrics on compact Clifford-Klein forms. The existence of some invariant geometric structures is an obstruction to compact Clifford-Klein forms, for example, none of para-Hermitian symmetric spaces admits a compact Clifford-Klein form [25].
Since the foundational works of Benoist, Labourie and T. Kobayashi [2],[3],[24] a lot of work has been done in the whole area. Since there are several excellent surveys of the whole topic (see [23],[25]) we prefer to refer to these works instead of a detailed exposition.
@@
In the context of this work, one should mention that there are many non-existence results for compact Clifford-Klein forms (not necessarily standard). We have already mentioned [6]. Topological obstructions to compact Clifford-Klein forms were found in [30],[31], [36] together with examples. A procedure of constructing homogeneous spaces with no compact Clifford-Klein forms was obtained in [7]. On the positive side, all compact standard Clifford-Klein of symmetric spaces were classified in [35]. The existence of non-standard compact Clifford-Klein forms is a deep mathematical problem of a different nature. For instance there are examples of (non-compact) Clifford-Klein forms which are in some sense strongly non-standard: in [9] one can find examples of homogeneous spaces $G/H$ of reductive type which admit proper actions of discrete subgroups of $G$ isomorphic to co-compact lattices in $O(n,1),$ $n=2,3,4$ but do not admit proper actions of any non-compact semisimple subgroups of $G.$ We also refer to a very recent work [29].
Finally, let us mention that we always consider the problem of standard Clifford-Klein forms under the assumption that $\mathfrak{h}$ and $\mathfrak{l}$ are reductive. The problem of existence of compact Clifford-Klein forms is meaningful also for other types of subgroups of $G$, for example for the solvable ones. There are no such forms for nilpotent $N\subset G$ [2] and for solvable $\Gamma\subset G$ under rather strong restrictions on $H$ [5].
1.2 Methods of proof of the main results
The proofs in this article are rather long and sometimes technical, therefore we feel it might be useful for the reader to have some overview first. Note that by [24] when looking for a standard compact Clifford-Klein form determined by a triple $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ one encounters with two somewhat different tasks:
  1. 1.
    one needs to know when a subgroup $L$ acts properly on $G/H$, and
  2. 2.
    when the double coset space $\Gamma\backslash G/H$ is compact.
The criteria of properness were found in [2] and [24]. They are expressed in Lie algebra terms and depend on the knowledge of the embeddings $\mathfrak{h}\hookrightarrow\mathfrak{g}$ and $\mathfrak{l}\hookrightarrow\mathfrak{g}$. This is one of the main difficulties if one wants to solve the general problem, since such knowledge may be not available. The second difficulty is to learn whether the given proper action is co-compact. According to [24] one needs to compare the non-compact dimensions of $\mathfrak{g},\mathfrak{h}$ and $\mathfrak{l}$ and prove the equality $d(\mathfrak{g})=d(\mathfrak{h})+d(\mathfrak{l})$ (still it is not an easy task, although the calculation does not depend on the embedding). Our methods fully use the results [21],[22], [24], but:
  1. 1.
    we derive a new property of the Iwasawa decompositions of $\mathfrak{g}$,$\mathfrak{h}$ and $\mathfrak{l}$ of any standard triple $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$. This property (called well-embeddedness ) is formulated in Lemma 1 and proved in [8] with the use of [1],
  2. 2.
    using Lemma 1 we derive some properties of the real root space decompositions of $\mathfrak{g},\mathfrak{h}$ and $\mathfrak{l}$ and their relations with root decompositions of the complexifications $\mathfrak{g}^{c}$,$\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ under the assumption that $\mathfrak{g}\not=\mathfrak{h}+\mathfrak{l}$ (this result was proved in [8], we repeat its formulation here, Theorem 4),
  3. 3.
    assuming now that the embedding $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$ is regular, we derive a contradiction between Theorem 4 and the regularity (Theorem 7),
  4. 4.
    in the non-regular case, we use Dynkin’s classification of maximal subalgebras of simple classical Lie algebras and under the assumptions that $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ are both simple and non-regular, we are able to derive the inequality $d(\mathfrak{g})>d(\mathfrak{h})+d(\mathfrak{l})$ which completes the whole proof.
The following remark seems to be in order. Our proofs essentially use the classification results from variuous sources: classical Dynkin’s papers [13],[14], tables of fundamental irreducible representations of simple Lie algebras [10],[11], tables of real forms of simple complex Lie algebras, of non-compact dimensions of simple Lie algebras and tables of maximal semisimple subalgebras in simple Lie algebras [33]. Parts of these used in this article are reproduced as Tables 2-6
Remark 1.
We were informed by the reviewer of our paper [8] that S. Mehdi and M. Olbrich announced in [28] that there are no compact standard Clifford-Klein forms of simple absolutely simple Lie groups $G$ and non-compact $H,L$ except those whose Lie algebras $\mathfrak{g},\mathfrak{h}$ and $\mathfrak{l}$ are contained in Table 1. The proof has not been published yet.
Acknowledgment. We thank Yves Benoist, Pralay Chatterjee, Toshiyuki Kobayashi, Yosuke Morita and T. N. Venkataramana for valuable discussions during the meeting "Zariski dense subgroups, number theory and geometric applications" held at ICTS, Bangalore in January 2024. Also, some parts of this work were completed during the second author research stay at the Institute of Mathematical Sciences in Chennai. His special thanks go to the Institute for a wonderful research atmosphere and hospitality.
2 Preliminaries
2.1 Notation
@@
Our basic references are [16],[17] and [33]. Assume that $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ is a standard triple, which means that $\mathfrak{g}$ is absolutely simple, $\mathfrak{h},\mathfrak{l}$ are reductive subalgebras of $\mathfrak{g}$ and $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ induces a compact standard Clifford-Klein form. Throughout we work with root systems of complex semisimple Lie algebras and with real root systems of their real forms. Also, we use some relations between them, which we reproduce here in order to fix notation (see [33], Section 4.1). We may assume that there exists a Cartan involution $\theta$ of $\mathfrak{g}$ such that $\theta(\mathfrak{h})=\mathfrak{h},$ $\theta(\mathfrak{l})=\mathfrak{l}.$ Hence the restriction of $\theta$ is a Cartan involution of $\mathfrak{h}$ and $\mathfrak{l}.$ We obtain the following Cartan decompositions
@@
$\mathfrak{g}^{c}=\mathfrak{j}^{c}+\sum_{\alpha\in\Delta}\mathfrak{g}_{\alpha}.$
The Lie algebra $\hat{\mathfrak{j}}:=i\mathfrak{t}+\mathfrak{a}$ is a real form of $\mathfrak{j}^{c}$ and
$\hat{\mathfrak{j}}=\{A\in\mathfrak{j}^{c}\ |\ \alpha(A)\in\mathbb{R}\ \textrm{for any}\ \alpha\in\Delta\}.$
The weights of the adjoint representation of $\mathfrak{a}$ constitute a root system $\Sigma\subset\mathfrak{a}^{*}$ (not necessarily reduced). There is a natural map $pr:\Delta\rightarrow\mathfrak{a}^{*}$ given by $\alpha\rightarrow\alpha|_{\mathfrak{a}}$. It is easy to see that $pr(\Delta)=\Sigma\cup\{0\}.$ Therefore, we will call $pr(\alpha)=\alpha|_{\mathfrak{a}}\in\mathfrak{a}^{*}$ (if non-zero) the restricted roots. Thus, $\Sigma=\{\alpha|_{\mathfrak{a}}\ |\ \alpha\in\Delta\}\setminus\{0\},$
$\mathfrak{g}_{\beta}:=\sum_{\begin{subarray}{c}\alpha\in\Delta\\ \alpha|_{\mathfrak{a}}=\beta\end{subarray}}\mathfrak{g}_{\alpha}\cap\ \mathfrak{g}.$
The reader should keep in mind, that we use the following convention. We use the same notation for complex roots in $\Delta$ and real (restricted) roots in $\Sigma$. The meaning of these will be always clear either from the context, or explicitly stated.
@@
$\mathfrak{g}=\mathfrak{m}_{0}+\mathfrak{a}+\sum_{\alpha\in\Sigma}\mathfrak{g}_{\alpha},$
where $\mathfrak{m}_{0}=\mathfrak{z}_{\mathfrak{g}}(\mathfrak{a})$. Let $B$ be the Killing form of $\mathfrak{g}.$ It extends onto $\mathfrak{g}^{c}$ as a Killing form of $\mathfrak{g}^{c}$ and we denote it by the same letter. Consider the compact real form $\mathfrak{g}_{u}:=\mathfrak{k}+i\mathfrak{p},$ of $\mathfrak{g}^{c},$ and denote by $\tau$ the complex conjugation in $\mathfrak{g}^{c}$ with respect to $\mathfrak{g}_{u}.$ Also we denote by $\sigma$ the complex conjugation of $\mathfrak{g}^{c}$ with respect to $\mathfrak{g}$. It is well known ([34], p. 228) that there is a Hermitian positive-definite form on $\mathfrak{g}^{c}$ given by the formula $B_{\tau}(X,Y)=-B(X,\tau(Y)).$ Choose a set of positive restricted roots $\Sigma^{+}\subset\Sigma.$ There is a system of positive roots $\Delta^{+}\subset\Delta$ such that
$\Sigma^{+}=\{\alpha|_{\mathfrak{a}}\ |\ \alpha\in\Delta^{+}\}-\{0\}.$
@@
as well as $\mathfrak{n}_{h}^{-},\mathfrak{n}_{l}^{-}$ for $\mathfrak{h}$ and $\mathfrak{l}.$ Since $\theta(\mathfrak{g}_{\alpha})=\mathfrak{g}_{-\alpha}$ for all $\alpha\in\Sigma$ ([16], Proposition 5.9) we get $\theta(\mathfrak{n})=\mathfrak{n}^{-}$. We will denote by $\Sigma_{h}$, $\Sigma_{l}$ the real root systems of $\mathfrak{h}$ and $\mathfrak{l}$, respectively, and consider $\mathfrak{n}_{h}$ and $\mathfrak{n}_{l}$ corresponding to $\Sigma_{h}$ and $\Sigma_{l}$.